User Tools

Site Tools


Sunday 5th April 16:00 CEST

  • second GRB observation - 090404 - this time first image (off position) at 6 seconds past GRB, first on position at 182 sec post GRB. This time without detection.

Thursday 26th March 13:30 CET

  • modelling is working - the telescope is autonomously building new model

Wednesday 25th March 22:00 NZDT

  • discovered problems with focusing - it does not hold on place due to bigger then expected weight of camera-filter wheel stack
  • continued autonomous operations without any problem
  • changed storage path for all-sky images

Tuesday 24th March 08:00 NZDT

  • first fully autonomous night without any presence runs as expected
  • problems with modelling due to wrong focus of the camera

Monday 16th March 12:00 NZDT

  • roof fixed (washers put back to expected position), UPS closing tests with telescope running performed - see this video.
  • Petr left Blenheim

Sunday 15th March 22:00 NZDT

  • make better focusing interface
  • telescope was focused at evening through open half of the roof
  • roof software was updated, error from Saturday fixed

Saturday 14th March 17:00 NZDT

  • roof stress tests showed weak point - wrong opening sequence can be achieved in some extreme circumstances. There is small damage on right half. We expect Cuddons will come on Monday, asset damage and put it back in shape.
  • we were able to close roof using some minor aid. Chain on sprocket wheels sometimes jump during closing.
  • preliminary observations showed no structure damage - washers holding transmission bearings were responsible for loosing energy accumulated in mechanism, and they are free a bit.

Thursday 12th March 23:20 NZDT

  • tested UPS closing - works as expected, but still I would like to test it with telescope moving
  • first truly remote opening - all looks good
  • tidy up cables
  • bought relay for Apogee remote control and 3 way switch (full manual, manual, automatic)

Thursday 12th March 01:00 NZDT

  • more cabling tidying - cables put to rack, rack fixed,..
  • soldered cable for cloud sensor - that takes at least 3 hours
  • Zelio upgrade tested and works as expected
  • after some showers it was finally clear during night, telescope works now in fully autonomous mode
  • UPS is able to handle operation of both halves of the roof with significant margin - that is better as expected

Wednesday 11th March 01:00 NZDT

  • roof opens and closes both halves simultaneously - with a small glitch, but this is already fixed and will be hopefully fully tested tomorrow
  • detection of emergency situations (e.g. timeouts) added to roof software and works as expected
  • auto-activation sequence (switch from standby to auto) works as expected
  • we hit overheat of the motors, both halves fall ~ 30 cm to opening supports (it looks that when motor overheats, clutch is released as well). No damage occurs and all motors operates properly after cold-down period
  • telescope driver fixed and operates almost perfectly
  • images are written to database
  • cloudsensor setpoints set (to 9 and 10), cloud sensor now have control over weather decisions
  • various updates and small fixes among code
  • fixed IP ordered, most probably that is why ADSL stop working properly. I will ask Phil tomorrow.

Tuesday 10th March 03:00 NZDT

  • Downloading USNO-1B, backup copy send on HDD by courier
  • Fixed telescope driver
  • Fixed roof controller issues with opening
  • Started upgrade of roof controller (Zelio logic)

Monday 9th March 10:00 NZDT

  • Observatory closed at morning - good
  • Telescope drive fails again - bad

Sunday 8th March 22:00 NZDT

  • Petr arrived at Blenheim
  • Telescope fixed by restarting control cabinet
  • rts2.ini updated with proper longitude and latitude
  • Observatory left in unattended mode

Saturday 7th March 18:00 NZDT

  • Andor rain protector fitted and telescope counterbalanced
  • Bottom rain protector fitted
  • Antonio leaves to Chile

Friday 6th March 18:00 NZDT

  • Dome mechanics fixed
  • Antonio arrives back at Blenheim

Wednesday 4th March 18:00 NZDT

  • Petr moves to Brisbane
  • Dome software bug finally discovered and fixed

Tuesday 3rd March 18:00 NZDT

  • Dome weather tests - still show minor issue after opening twice
  • Successful UPS opening/closing tests
  • Tidy-up cabling, make and use holes for all cables, so the place looks nice and tidy
  • Tested telescope movement at full speed

Monday 2nd March 18:00 NZDT

  • Dome tests
  • Finished new cabling for cloud and rain sensors

Saturday 1st March 18:00 NZDT

  • Control room finished (in one day)
  • Server rack moved to control room
  • Final tests of telescope by Astelco
  • Andor CCD connected
  • Telescope first light at final optical configuration
  • All sky CCD worked quite well

Friday 27th February 12:00 NZDT

  • Dome cabling complete
  • Dome tested with Zelio, one minor issue fixed
  • Emergency stops installed and readied
  • Everything ready for opening ceremony

Thrusday 26th February 24:00 NZDT

  • Telescope assembled
  • Dome cabling finished to level that opening of domes is possible through relays
  • Telescope tested, with preliminary model created

Wednesday 25th February 18:00 NZDT

  • Both halfs of dome tested, looks much better then before
  • Alberto arrived
  • Welding of telescope support commence

Tuesday 24th February 18:00 NZDT

  • Dome modifications on one half almost done, tests expected tomorrow
  • Astelco started telescope assembly
  • Server rack moved a bit in preparation of its relocation to the dome

Monday 23rd February 18:00 NZDT

  • Rest of telescopes parts arrived
  • Telescope interface ready at Blenheim workshop
  • FLI ↔ Andor interface being manufactured
  • Dome redesigned, some work on redesign done. First tests are expected on Tuesday afternoon

Friday 20th February 18:30 NZDT

  • Andor images successfully retrieved - the issue was system having more then 3GB RAM. We are able to get images when OS uses 3GB. There is a request to Andor to fix this issue
  • Design of roof transmission is finished, we expect some testing tomorrow. So far nobody from the company showed, but they phoned they will arrive with transmission on Saturday

Thursday 19th February 24:00 NZDT

  • roof will be hopefully fixed by added 1:3 - 1:4 sprocket wheel - chain gearbox. On Friday we shall have detailed design
  • we do not know how will roof behaves during strong winds. If we can found some aerodynamic expert, that can provide us help
  • Andor is working, but giving 0s at CCD output. My estimate is that this is software problem. I will run more tests on Friday
  • Cloud sensors show erratic behaviour (works now). I will try to contact author to give hints where it can be. We suspect RS232 connector short
  • ssh access to b3-main (port 2222) and b3-allsky (port 2223) works; you will need to ask for password
  • Florentino complain about heat at sensor rack. ipmitool, which were installed, does not shows any significant heat - but I will keep an eye on that
  • new Zelio arrived and was tested - works fine
  • Spanish ambassador called Bill and confirmed his attendance at opening (with a box of Spanish wines)

Wednesday 18th February 18:00 NZDT

  • second half of the dome installed
  • Interned connection to Bootes 3 computers works
  • weather mast installed - cloud sensor, Weather, Outside webcam, rain sensor
  • UPS tested

Tuesday 17th February 19:00 NZDT

  • Dome opened a bit by motors
  • Second half of the roof almost finish, ready for tomorrow placement on top
  • Weather station mast welded
  • Davis weather and cloud sensors left on overnight for tests, logging data to computers
  • ADSL connection fixed
  • AXIS camera confirmed working
  • Bought cables for network (AXIS, Zelio, ..)
  • Andor camera shows problems when exposing (kernel panic), I will look for solutions tomorrow (expected is that driver change will fix it)

Tuesday 17th February 09:30 NZDT

  • Andor power supply connector arrived
  • ordered new Zelio logic through local dealer (~ 700 NZD) and 24 V power supply (~ 160 NZD)
  • both servers successfully ran overnight

Monday 16th February 18:00 NZDT

  • Antonio arrived
  • one half of the roof was lifted and installed, Florentino is working to make motors operating
  • second half of the roof final assembly begins, one plate (out ot 6) is on place
  • bought Ethernet cables (~50 NZD) and ordered connector for additional Andor power (SATA→IDE interface, ~14 NZD)
  • successfully tested Apogee CCD and cloud sensor
  • servers preliminary integrated in rack for testing
  • mount arrives in a box, we are looking for interface to connect it to pier
  • it was discovered that we do not have power supply for Andor CCD cooling. CCD will be operational without it, Astelco should carry power supply

Sunday 15th February 14:00 NZDT

  • finished installation of the servers, going to test them with hardware
  • bought 1GB switch (187 NZD ~ 100 EUR)
  • yesterday we find out that the current ADSL Internet connection is way slow. Most probable culprit is missing splitter, that will be installed on Tuesday. Final Internet connection will be solved by own line, which will hopefully be installed sometime next week
  • also there is a problem with Zelio logic control box - the one bought is incompatible with network module. Temporal solution is to run roof without network control in manual mode, final solution will be in replacement of the Zelio control. The problem is pending confirmation from Schneider-Electrics.

Saturday 14th February 12:00 NZDT

  • dome is nearing completion
  • one half of the roof is complete, second half waits for plates to be mated to sites - that will take a few days. I expect that will be finished on Tuesday or Wednesday
  • Florentino is working on motors, hopefully they will be ready within few days
  • our servers are up and running, I am installing operation system and RTS2 on them
  • looking for options how to get network connected to the observatory
obs/bootes_3_constr.txt · Last modified: 2009/04/05 00:00 (external edit)